Jarvene Corporation uses the FIFO method in its process costing system. The following data are for the most recent month of operations in one of the company’s processing departments:

Units in beginning inventory 420
Units started into production 4,290
Units in ending inventory 320
Units transferred to the next department 4,390
Materials Conversion
Percentage completion of beginning inventory 80 % 20 %
Percentage completion of ending inventory 70 % 40 %

The cost of beginning inventory according to the company’s costing system was $7,823 of which $4,851 was for materials and the remainder was for conversion cost. The costs added during the month amounted to $178,986. The costs per equivalent unit for the month were:

Materials Conversion
Cost per equivalent unit $18.00 $23.00

Required:

1. Compute the total cost per equivalent unit for the month.

2. Compute the equivalent units of material and conversion in the ending inventory.

3. Compute the equivalent units of material and conversion that were required to complete the beginning inventory.

4. Compute the number of units started and completed during the month.

5. Compute the cost of ending work in process inventory for materials, conversion, and in total for the month.

6. Compute the cost of the units transferred to the next department for materials, conversion, and in total for the month.
Answer #1

1) Total cost per equivalent unit = 18+23 = 41

2) Equivalent unit of ending inventory

Material Conversion
Equivalent unit in ending inventory 320*70% = 224 320*40% = 128

3) Cost of equivalent unit in beginning inventory completed

Material Conversion
Equivalent unit in beginning inventory 420*20% = 84 420*80% = 336

4) Cost of unit started and completed = 4390-420 = 3970

5) Cost of ending WIP

Material Conversion Total
Cost of ending WIP 224*18 = 4032 128*23 = 2944 6976

6) Cost of units transferred out

Material Conversion Total
Cost of unit transferred out (4054*18+4851) = 77823 (4306*23+2972) = 102010 179833